OpenStudy (anonymous):

What value of s makes the statement true? What value of s makes the statement true? (–7 + s) + (–14) = (12 + (–7)) + (–14) A. –30 B. –12 C. –7 D. 12

OpenStudy (agl202):

What do u think?

OpenStudy (anonymous):

I am thinking

OpenStudy (agl202):

-7+s - 14 = 12 - 7 - 14 -21 + s = -9 s = 12 D. is correct answer! did u get it?